ASSESSMENT




Perform as indicated




In numbers 1-5, find the critical valuets) and rejection remonts) for the type of z-test with level of significance a Include a graph with your answer




1 Left -tailed test, a=0.03




2 Right-tailed test = 0.05




3 Two-tailed test, a = 0.02




4 Two-tailed test, a = 0.10 5 Left-tailed test, a=0,09




In numbers 6-9, state whether each standardized test statistic z allows you to reject the null hypodesis Explain your reasoning




6. z=-1301 7 z=1203




8 2 1.280 9 2=1.286




10 A fast food restaurant estimates that the mean sodium content in one of its breakfast sandwiches is no




more than 920 milligrams A random sample of 44 breakfast sandwiches has a mean sodium content of 925 milligrams Aanume the population standard deviation is 15 milligrams At a=0.10, do you have enough evidence to reject the restaurant's claim?

1
Expert's answer
2022-05-27T14:39:13-0400

1.


"z\le-1.8808"



2.


"z\ge1.6449"


3.


"|z|\ge2.3263"



4.


"|z|\ge1.6449"



5.


"z\le -1.3408"


Let we have left-tailed and "z_c=-1.2816"

6.

Reject "H_0" because "z< \u22121.2816."


7. Fail to reject "H_0" because "z > \u22121.2816."


8. Fail to reject "H_0" because "z > \u22121.2816."


9. Fail to reject "H_0" because "z > \u22121.2816."


10.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le920"

"H_1:\mu>920"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.10," and the critical value for a right-tailed test is "z_c = 1.2816."

The rejection region for this right-tailed test is "R = \{z:z>1.2816\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{925-920}{15/\sqrt{44}}\approx2.211"

Since it is observed that "z=2.211>1.2816=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(z>2.211)=0.027036," and since "p= 0.027036<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 920, at the "\alpha = 0.10" significance level.